OpenStudy (anonymous):

Which equation would you use to solve the following word problem? Sasha's mom decided to get her a cell phone for her birthday. Cell Plus has a plan that costs $24.95 per month plus an additional $0.06 per minute. Cell Best has a plan that costs $30.95 per month plus an additional $0.05 per minute. How many minutes can Sasha talk and have the same cost each month? I NEED AN ANSWER FOR AN ONLINE TEST PLEASE HELP NOW!!!! :'(

OpenStudy (anonymous):

A) 24.95 + 0.06m = 30.95 + 0.05m B) 0.11m = 6.00 C) 24.95 + 0.06m = 31.00 D) 30.95 + 0.05m = 25.01

OpenStudy (anonymous):

@JakeV8 please help!

OpenStudy (anonymous):

Right answer is A

OpenStudy (anonymous):

you just write out the expression for the cost of each plan, and then set the 2 expressions to be = because the problem asked how many minutes would make the total cost the same. Answer a shows Cell Plus plan on the left and Cell Best on the right

OpenStudy (anonymous):

if you actually have to solve for the number of minutes, simplify the equation in answer a, and then solve for m, the # of minutes that makes the two plans equal cost.

OpenStudy (anonymous):

Thx @JakeV8 and @ktnguyen1